OpenStudy (wannabegurl):

What's 25/36 as a exponent ? Or whole #

OpenStudy (abdullahm):

\(\sf\Large \frac{a^2}{b^2} = \left(\frac{a}{b}\right)^2\)

OpenStudy (abdullahm):

So... \(\sf\large\frac{25}{36} = \frac{5^2}{6^2} = ?\)

OpenStudy (wannabegurl):

I don't understand ...

OpenStudy (wannabegurl):

omg , thanks !

OpenStudy (abdullahm):

np